Problem 5.102 Partially correct answer. Your answer is partially correct. Try again. The percentage of people given an antirheumatoid who suffer severe, moderate, or minor side effects are 13, 20 and 67%, respectively. Assume that people react independently and that 20 people are given the medicine. Determine the following. Round your answers to four decimal places (e.g. 98.7654). (a) The probability 2, 4, and 14 people will suffer severe, moderate, or minor side effects, respectively. Probability = Entry field with correct answer (b) The probability that no one will suffer severe side effects. Probability = Entry field with correct answer (c) The mean and variance of the number of people who will suffer severe side effects. Mean = Entry field with correct answer Variance =
Answer #1

a)

The probability 2, 4, and 14 people will suffer severe, moderate, or minor side effects

= =0.0577

b)

probability that no one will suffer severe side effects =(0.87)20 =0.0617

c)

mean =np=20*0.13=2.6

Variance = np(1-p)=2.262